### v1

# Robotic Learning for Informative Path Planning

Abstract
[…]
AIPP is important to many robotics applications, enabling mobile robots to efficiently collect useful data about initially unknown environments. In addition, learning-based methods are increasingly used in robotics to enhance adaptability, versatility, and robustness across diverse and complex tasks. Our survey explores research on applying robotic learning to adaptive informative path planning (AIPP), bridging the gap between these two research fields. We begin by providing a unified mathematical framework for general AIPP problems. Next, we establish two complementary taxonomies of current work from the perspectives of (i) learning algorithms and (ii) robotic applications. We explore synergies, recent trends, and highlight the benefits of learning-based methods in AIPP frameworks. Finally, we discuss key challenges and promising future directions to enable more generally applicable and robust robotic data-gathering systems through learning. We provide a comprehensive catalogue of papers reviewed in our survey, including publicly available repositories, to facilitate future studies in the field.
keywords:
Informative path planning, Robot learning, Active learning
[…]
figurec
2D
two-dimensional
3D
three-dimensional
AHRS
attitude and heading reference system
AUV
autonomous underwater vehicle
COMA
counterfactual multi-agent policy gradients
CPP
Chinese Postman Problem
DoF
degree of freedom
DVL
Doppler velocity log
FSM
finite state machine
IMU
inertial measurement unit
LBL
Long Baseline
MCM
mine countermeasures
MDP
Markov decision process
MCTS
Monte Carlo Tree Search
POMDP
partially observable Markov decision process
PRM
Probabilistic Roadmap
RL
reinforcement learning
ROI
region of interest
ROS
Robot Operating System
ROV
remotely operated vehicle
SLAM
simultaneous localization and mapping
SSE
sum of squared errors
STOMP
Stochastic Trajectory Optimization for Motion Planning
TRN
Terrain-Relative Navigation
UAV
unmanned aerial vehicle
UGV
unmanned ground vehicle
USBL
Ultra-Short Baseline
IPP
informative path planning
AIPP
adaptive informative path planning
FoV
field of view
CDF
cumulative distribution function
ML
maximum likelihood
RMSE
Root Mean Squared Error
MLL
Mean Log Loss
GP
Gaussian Process
KF
Kalman Filter
IP
Interior Point
BO
Bayesian Optimization
SE
squared exponential
UI
uncertain input
MCL
Monte Carlo Localisation
AMCL
Adaptive Monte Carlo Localisation
SSIM
structural similarity index measure
PSNR
peak signal-to-noise ratio
LPIPS
learned perceptual image patch similarity
MAE
Mean Absolute Error
RMSE
root mean squared error
AUSE
Area Under the Sparsification Error curve
mIoU
mean Intersection-over-Union
CMA-ES
covariance matrix adaptation evolution strategy
RRT
rapidly exploring random tree
RIG
rapidly exploring information gathering
AL
active learning
MLP
multi-layer perceptron
LSTM
long short-term memory
PPO
proximal policy optimization
CNN
convolutional neural network
IL
imitation learning
GPS
Global Positioning System
[…]
[label1]organization=MAVLab, Faculty of Aerospace Engineering, TU Delft,addressline=Kluyverweg 1, city=Delft, postcode=2629 HS, country=The Netherlands
[…]
[label2]organization=Cluster of Excellence PhenoRob, Institute of Geodesy and Geoinformation, University of Bonn,addressline=Niebuhrstr. 1A, city=Bonn, postcode=53113, country=Germany
[…]
[label3]organization=Stanford Intelligent Systems Laboratory, Department of Aeronautics and Astronautics, Stanford University,addressline=496 Lomita Mall, city=Stanford, postcode=94305, state=California, country=United States of America
1 Introduction
The field of robotics has witnessed remarkable advancements in recent years, fueled by the growing need for automation and the increasing complexity of tasks required in various application domains. One of the key challenges in robotics is AIPP, which involves planning a trajectory for an autonomous robot to follow that maximizes the information acquired about an unknown environment while simultaneously respecting its resource constraints. This problem is of critical importance in applications such as environmental monitoring, exploration, search and rescue, and inspection across ground, aerial, and aquatic domains (Rayas Fernández et al., 2022; Cao et al., 2023a; Popović et al., 2020; Denniston et al., 2023a; Rückin et al., 2022a; Viseras and Garcia, 2019; Choudhury et al., 2020).
Despite its importance, AIPP is a challenging problem due to the inherent complexity of modeling and predicting new information in the environment, as well as the need to balance the exploration of unknown areas with the exploitation of newly acquired data (Singh et al., 2009). Additionally, noisy sensor measurements introduce uncertainties into the data acquisition process, while actuation uncertainty can lead to deviations from planned trajectories. Moreover, real-world environments are often highly dynamic, which further complicates the modeling process and makes prediction challenging. Decisions must be made in a sequential manner to allow for continual refinement as more information becomes available during a mission.
Conventional approaches such as static pre-computed paths tend to fail on AIPP problems because they often rely on strong assumptions about the environment and cannot adapt to uncertainties or changes online (Galceran and Carreras, 2013). Additionally, these approaches scale poorly to large, complex environments, and may not effectively account for the constraints and capabilities of the robot itself. These conventional solutions applied to the AIPP problem have been limited by computational complexity, lack of adaptability, and an inability to generalize across diverse environments and problems (Tan et al., 2021). The growing popularity of learning-based methods has led to a renewed focus on applying these new techniques to AIPP, offering the promise of more flexible, adaptive, and scalable solutions.
In this survey paper, we present a new perspective on AIPP by exploring the potential of emerging robot learning techniques in addressing the challenges of AIPP. Different learning-based methods allow for constructing path planning algorithms that can naturally account for the inherent uncertainties and dynamic changes within a given environment. This adaptability facilitates more robust and efficient operations in complex environments because the robot can optimize its path based on its evolving understanding of its surroundings. This not only improves task performance, but also enhances resilience and versatility since the robot can react more effectively to new information about its environment.
Our review includes different aspects of learning: supervised learning, reinforcement learning, imitation learning, and active learning. We focus on how these techniques can be used for AIPP in robotics. Furthermore, we discuss relevant application domains, such as environmental monitoring, exploration and search, semantic scene understanding, and active simultaneous localization and mapping (SLAM).
While there have been previous surveys addressing aspects of AIPP (Bai et al., 2021; Maboudi et al., 2023; Lluvia et al., 2021; Sung et al., 2023; Taylor et al., 2021; Aniceto and Vivaldini, 2022), our work distinguishes itself by specifically focusing on the intersection of learning-based methods and AIPP. Previous literature reviews have focused on specific subfields of robotic learning. For instance, Taylor et al. (2021) focus on active learning in robotics, highlighting methods suitable for the demands of embodied learning systems. Similarly, Argall et al. (2009) address imitation learning, where robots develop policies from example state-to-action mappings, while Garaffa et al. (2023) investigate reinforcement learning techniques that devise unknown environment exploTration strategies for single and multi-robot exploration. Lauri et al. (2022) exclusively examine the application of Partially Observable Markov Decision Processes (POMDPs) in robotics, and Mukherjee et al. (2022) analyze robotic learning strategies for human-robot collaboration in industrial settings. Other surveys, such as those by Sung et al. (2023), Aniceto and Vivaldini (2022), and Bai et al. (2021) explore different different applications of AIPP, such as environmental monitoring tasks or general path planning techniques (Bai et al., 2021; Galceran and Carreras, 2013; Chen, 2011; Tan et al., 2021; Maboudi et al., 2023). However, these works do not underscore the applicability of learning in adaptive scenarios.
The main objectives of this survey paper are to:
  1. 1.
    provide a comprehensive understanding and taxonomy of the current state-of-the-art in learning-based methods for AIPP as well as their applications;
  2. 2.
    introduce a standardized mathematical framework for AIPP problems, which provides a unified foundation for understanding and comparing various learning-based methods and application domains;
  3. 3.
    identify potential avenues for future research by highlighting the limitations of current approaches.
We have cataloged the papers in our survey at https://dmar-bonn.github.io/aipp-survey/. In particular, we highlight papers that have open-source implementations available.
2 Mathematical Formulation
We study the general problem of adaptive informative path planning (AIPP). The goal is to find an optimal action sequence $\psi^{*}=(a_{1},\ldots,a_{N})$ of $N$ robot actions $a_{i}\in\mathcal{A},i\in\{1,\ldots,N\},$ where actions may be, e.g., acceleration commands, next robot poses, using a specific sensor, or returning to a charging station. The action sequence $\psi^{*}$ maximizes an information-theoretic criterion $I(\cdot)$:
$\psi^{*}=\operatorname*{argmax}_{\psi\in\Psi}\,I(\psi),\,\text{s.t. }C(\psi)\leq B\,,$ (1)
where $\Psi$ represents the set of all possible action sequences, the cost function $C:\Psi\to\mathbb{R}$ maps an action sequence to its associated execution cost, $B\in\mathbb{R}$ is the robot’s budget limit, e.g., time or energy, and $I:\Psi\to\mathbb{R}$ is the information criterion, computed from the new sensor measurements obtained by executing the actions $\psi$.
Our problem setup considers a robot gathering sensor measurements in an initially or partially unknown environment. In this environment $\xi\subset\mathbb{R}^{D}$, AIPP requires online replanning during a mission to adaptively focus on areas of interest as they are discovered. Each point $\mathbf{x}\in\xi$, e.g., a $D$-dimensional pose within the environment, is characterized by its distinctive features, represented as $F(\mathbf{x})\in\mathcal{F}$. Here, $\mathcal{F}$ is a feature space, which could include characteristics like the semantic class, temperature, or radiation level. These features may fluctuate during the mission, as defined by the feature mapping function $F:\xi\to\mathcal{F}$. The notation used to describe our AIPP formulation is summarized in Table 1.
Areas of interest within the environment the robot operates in are characterized by their features, which indicate interesting phenomena as defined by the mission objectives. Examples of mission objectives include identifying victims in a post-disaster scenario or locating regions of high radiation levels in a nuclear facility monitoring task. This concept of adaptivity in an optimal action sequence $\psi^{*}$ is reflected in the specific definition of the information-theoretic criterion $I$ during the optimization problem presented in Eq. 1.
Importantly, the point features $F(\mathbf{x})$, such as temperature, may change over time. Given that the initial environment is unknown, the robot’s understanding of $F(\mathbf{x})$ may need to be updated during the mission. This could result in updating beliefs regarding interesting regions over time as new sensor measurements are received. Thus, to ensure that the robot’s behavior is adaptively informed, we may re-solve the constrained optimization problem in Eq. 1 in a computationally efficient manner during a mission.
Symbol Meaning Examples
$a_{i}\in\mathbb{A}$ robot action next robot position or pose, sensing behavior
$\psi\in\Psi$ sequence of consecutive actions
$\mathbf{x}\in\xi$ areas/locations in an environment surfaces, patches of terrain
$I:\Psi\to\mathbb{R}$ information-theoretic criterion map entropy, mutual information
$F:\xi\to\mathcal{F}$ feature mapping function semantic class, spatial occupancy, temperature, radiation level
Table 1: Notation associated with the adaptive informative path planning (AIPP) problem.
3 Background
Figure 1: System diagram showing the key elements of a general AIPP framework. During a mission, a map $\tilde{F}$ of the robot’s environment $\xi$ is built using measurements $z$ extracted from a sensor data stream. An AIPP algorithm leverages the map data to find the next action sequences $\psi$ for the robot to maximize the information value ${I}(\psi)$, i.e. utility, of future measurements. The next action sequence $\psi$ is executed by the robot, allowing for subsequent map updates in a closed-loop manner.
The generic structure of a system for AIPP is shown in Fig. 1. In an active sensing context, the framework processes raw data from a robot sensor to obtain measurements used to build a map of the robot’s environment. During a mission, the AIPP planning algorithm uses the environment maps built online to optimize future action sequences for maximum gain in an information metric as defined by the utility function. This section describes methods for environment mapping, evaluation metrics, and standard benchmarks commonly used in AIPP frameworks.
3.1 Mapping
The AIPP problem assumes the environment $\xi$ to be a priori (partially) unknown. Thus, the true feature mapping $F:\xi\to\mathcal{F}$ and associated true regions of interest are a priori (partially) unknown as well. This makes the original AIPP problem stated in Eq. 1 not only computationally challenging but also impossible to solve directly as exact knowledge about interesting regions requires ground truth access to the feature mapping $F$. Instead of solving Eq. 1 directly, we model a stochastic process $\tilde{F}$ over all possible feature mapping functions. At each mission time step $t$, the robot’s belief about the true feature mapping $F$, i.e., the probability measure associated with $\tilde{F}$ conditioned on the action-observation history, is updated by:
$p(F\mid z_{1},\ldots,z_{t},a_{1},\ldots,a_{t})\,,$ (2)
where $\{z_{1},\ldots,z_{t}\}$ and $\{a_{1},\ldots,a_{t}\}$ are all previously collected measurements and executed actions, respectively. The belief about $\tilde{F}$ is used to approximate the true information criterion $I$ in Eq. 1.
To approximate the AIPP problem in Eq. 1 as closely as possible, the robot’s belief about $\tilde{F}$ in Eq. 2 may be updated online during a mission as new sensor data $z_{t}$ arrives at time step $t$ after executing a sensing action $a_{t}$. Note that the measurements $z_{t}$ at time step $t$ may be uncertain or noisy. The measurements $z_{t}$ are assumed to be sampled from $p(z\,|\,F)$ representing a probabilistic sensor model. Furthermore, there may be uncertainty in localization and actuation, leading to imperfect estimation of the robot pose as measurements are taken. In Sec. 4 and Sec. 5, we discuss in detail how these different sources of uncertainty can be accounted for in the context of learning-based methods.
To compute the belief $\tilde{F}$ following Eq. 2 based on the collection of stochastic measurements $\{z_{1},\ldots,z_{t}$}, various methods for environment mapping are commonly used. Table 2 provides an overview of commonly used mapping methods and corresponding metrics derived from them used to evaluate the performance of AIPP methods. We consider four broad categories in our taxonomy: occupancy grid mapping (Elfes, 1989), Gaussian processes (Rasmussen and Williams, 2006), graph-based methods, and implicit neural representations (Mildenhall et al., 2021). In general, the chosen mapping method for a given AIPP scenario depends on the environment, task, and available computing resources. For instance, Gaussian processes are applicable for problems where the to-be-mapped features $F(\mathbf{x})$ support continuous feature spaces $\mathcal{F}$ and the map supports continuous environment representations $\xi$, whereas occupancy grid mapping is suitable for discrete mapping tasks. In contrast, graph-based methods are typically used for active SLAM problems (Placed et al., 2023) where both environment mapping and robot localization are concurrent. Recently, implicit neural representations, such as NeRFs (Mildenhall et al., 2021) and pixelNeRFs (Yu et al., 2021), are gaining popularity for AIPP without relying on an explicit environment map. Due to their lightweight nature, these methods enable conserving computational and memory resources during planning in 3D reconstruction tasks where a high level of detail is required.
Evaluation metrics References
Occupancy grid Entropy (Popović et al., 2020; Westheider et al., 2023; Bai et al., 2017; Chen et al., 2020; Sukkar et al., 2019; Choudhury et al., 2018)
Mutual information (Lodel et al., 2022; Chen et al., 2019)
Root mean squared error (Popović et al., 2020)
Mean absolute error (Chen et al., 2021; Viseras et al., 2021)
Mean intersection-over-union (Georgakis et al., 2022a; Ramakrishnan et al., 2020)
F1-score (Westheider et al., 2023)
Accuracy (Schmid et al., 2023; Georgakis et al., 2022a; Ramakrishnan et al., 2020)
Precision/recall (Schmid et al., 2023)
Coverage (Schmid et al., 2023, 2022; Niroui et al., 2019; Cao et al., 2023b; Reinhart et al., 2020; Zeng et al., 2022; Shrestha et al., 2019; Chen et al., 2021; Zacchini et al., 2023; Liu et al., 2023; Vasquez-Gomez et al., 2021; Mendoza et al., 2020; Caley et al., 2019; Georgakis et al., 2022a; Ramakrishnan et al., 2020; Chen et al., 2022; Tao et al., 2023; Li et al., 2023; Gao and Zhang, 2022; Sukkar et al., 2019; Zwecher et al., 2022; Dai et al., 2022)
Observed surfaces (Hepp et al., 2018; Dhami et al., 2023; Pan et al., 2023a)
Landmark error (Choudhury et al., 2020)
Gaussian process Entropy (Denniston et al., 2023b)
Expected improvement (Denniston et al., 2023b; Song et al., 2015)
Probability of improvement (Denniston et al., 2023b)
Covariance matrix trace (Cao et al., 2023a; Rückin et al., 2022a; Popović et al., 2020, 2020; Ott et al., 2023; Yang et al., 2023a; Yanes Luis et al., 2023)
Mutual information (Hitz et al., 2017; Wei and Zheng, 2020)
Root mean squared error (Cao et al., 2023a; Rückin et al., 2022a; Rayas Fernández et al., 2022; Vivaldini et al., 2019; Popović et al., 2020, 2020; Hollinger and Sukhatme, 2014; Viseras and Garcia, 2019; Yanes Luis et al., 2023; Choi and Cielniak, 2021)
Mean log likelihood (Popović et al., 2020)
Distribution divergence (Cao et al., 2023c)
Mean uncertainty (Cao et al., 2023c; Hollinger and Sukhatme, 2014)
Upper confidence bound (Marchant et al., 2014)
Graph-based Landmark uncertainty (Chen et al., 2021, 2020)
Landmark error (Chen et al., 2020)
Localization uncertainty (Chen et al., 2020; Popović et al., 2020)
Localization error (Popović et al., 2020; Hüttenrauch et al., 2019)
Target uncertainty (Duecker et al., 2021; Tzes et al., 2023; Singh et al., 2009)
Detection rate (Best et al., 2019)
Path costs (Meliou et al., 2007)
Neural representation Peak signal-to-noise ratio (Jin et al., 2023; Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023; Sünderhauf et al., 2023; Pan et al., 2023b)
Structural similarity index measure (Jin et al., 2023; Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023; Pan et al., 2023b)
Learned perceptual image patch similarity (Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023)
Accuracy (Ran et al., 2022; Zhan et al., 2023; Yan et al., 2023)
Completion (Ran et al., 2022; Zhan et al., 2023; Yan et al., 2023)
Completion ratio (Ran et al., 2022; Yan et al., 2023)
F1-score (Zhan et al., 2023; Lee et al., 2022)
Chamfer distance (Zhan et al., 2023)
Precision (Lee et al., 2022)
Recall (Lee et al., 2022)
Surface coverage (Lee et al., 2022; Yan et al., 2023)
Table 2: Mapping methods and evaluation metrics used in AIPP.
3.2 Evaluation Metrics
The goal of AIPP is to adaptively acquire information about initially unknown environment features $F$ by collecting uncertain or noisy measurements $z_{t}$ during a mission, usually captured as a posterior map belief $\tilde{F}$ in Eq. 2. Because the robot mission goals are task- and application-dependent, e.g., searching for targets (Ott et al., 2023; Tzes et al., 2023; Cao et al., 2023c) or regions of high surface temperature (Rückin et al., 2022a; Meliou et al., 2007; Westheider et al., 2023), there is a lack of standard evaluation metrics used to quantify the performance of AIPP algorithms. Most works choose evaluation metrics closely related to their specified information criterion $I$ in Eq. 1 because they reflect or approximate the mission goal. Table 2 lists commonly used evaluation metrics. Although these metrics vary significantly in the AIPP community, most works consider quantitative measures based on the map representation to compute the belief $\tilde{F}$, as reflected by our taxonomy.
Applications aiming to collect information about discrete features $F$, such as spatial occupancy (Bai et al., 2017; Chen et al., 2020) or land cover classification Popović et al. (2020); Vivaldini et al. (2019); Rayas Fernández et al. (2022), commonly use the entropy of the belief $\tilde{F}$ as a measure for the remaining uncertainty of the environment features (Popović et al., 2020; Westheider et al., 2023; Bai et al., 2017; Chen et al., 2020; Sukkar et al., 2019; Choudhury et al., 2018). Lower entropy of $\tilde{F}$ indicates better AIPP performance. Following popular computer vision benchmarks Everingham et al. (2010); Cordts et al. (2016), some works evaluate the classification quality of $\tilde{F}$ by computing the mean Intersection-over-Union (mIoU(Georgakis et al., 2022a; Ramakrishnan et al., 2020), accuracy (Schmid et al., 2023; Georgakis et al., 2022a; Ramakrishnan et al., 2020), or F1-score (Westheider et al., 2023) of the maximum a posteriori estimate of the map belief $\tilde{F}$ given the ground truth features $F$.
Methods gathering information about continuous features $F$ in an environment, e.g., bacteria level in a lake (Hitz et al., 2017) or magnetic field strength (Viseras and Garcia, 2019), often use Gaussian process models Hitz et al. (2017); Cao et al. (2023a); Viseras and Garcia (2019); Ott et al. (2023) or some variant of Bayesian filtering Popović et al. (2020); Rückin et al. (2022a); Choudhury et al. (2020); Denniston et al. (2023a) as the map representations. Although Gaussian processes allow for computing the differential entropy of $\tilde{F}$ in closed-form (Williams and Rasmussen, 2006), it is common to use the covariance matrix trace to approximate the remaining uncertainty about the environment features and hence the information criterion $I$ due to its computational efficiency. Despite relaxed computational requirements, most works evaluate AIPP performance based on the covariance matrix trace during offline evaluation as well (Cao et al., 2023a; Rückin et al., 2022a; Popović et al., 2020, 2020; Ott et al., 2023; Yang et al., 2023a; Yanes Luis et al., 2023). To evaluate the reconstruction quality of $\tilde{F}$, the root mean squared error (RMSE) between the maximum a posteriori estimate of $\tilde{F}$ and the ground truth $F$ is usually reported Cao et al. (2023a); Rückin et al. (2022a); Rayas Fernández et al. (2022); Vivaldini et al. (2019); Popović et al. (2020, 2020); Hollinger and Sukhatme (2014); Viseras and Garcia (2019); Yanes Luis et al. (2023); Choi and Cielniak (2021).
Evaluation methods for implicit neural representations differ in that they only rely on a trained neural network to represent the entire environment. The metrics can be grouped based on two aspects: the quality of the rendered images and the quality of the geometry of the reconstructed surface. Image quality metrics, e.g., peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS), are computed based on a set of evenly distributed test images in the environment (Jin et al., 2023; Pan et al., 2022; Sünderhauf et al., 2023; Ran et al., 2022; Zhan et al., 2023; Pan et al., 2023b). To evaluate the reconstruction quality, the learned geometric representation is usually first post-processed, e.g., by using the marching cubes algorithm to extract a mesh. Then, a set of points is sampled on the mesh to compute various metrics, e.g., mesh accuracy or completion (Lee et al., 2022; Ran et al., 2022; Zhan et al., 2023).
Some methods do not assume the robot to have perfect ground truth localization information during missions but additionally account for localization uncertainty while building the map belief $\tilde{F}$. These works propagate localization uncertainty into the map belief update in Eq. 2 using graph-based SLAM systems (Chen et al., 2020, 2021; Popović et al., 2020; Tzes et al., 2023) and Gaussian process variants supporting input uncertainty (Popović et al., 2020). In addition to assessing the quality of the map, the quality of robot localization is also evaluated to determine whether the AIPP method can gather not only informative measurements to improve the environment map, but also measurements that improve localization to guarantee the construction of an accurate map. Some works compute the trajectory errors between the estimated and ground truth executed trajectory Chen et al. (2021). Other works track the localization uncertainty of the robot pose belief Popović et al. (2020); Chen et al. (2020), the localization error between the estimated and ground truth robot pose Popović et al. (2020); Hüttenrauch et al. (2019), or the landmark errors between estimated and ground truth landmark poses (Chen et al., 2020).
An alternative problem formulation considers using AIPP to actively improve the sensor model reasoning about measurements $z_{t}$. In Eq. 1, this is done by linking the information criterion $I$ to the value of new training data acquired during a mission. Such formulations generally do not aim to improve the belief of the map $\tilde{F}$ representing a physically quantifiable environmental phenomenon. Instead, works in this category target improving a neural network for semantic segmentation using RGB images (Blum et al., 2019; Zurbrügg et al., 2022; Rückin et al., 2023, 2024) or object detection (Chaplot et al., 2021). AIPP performance is evaluated using a held-out test set to quantify the neural network prediction performance Everingham et al. (2010); Cordts et al. (2016). Commonly used metrics include mIoU Blum et al. (2019); Zurbrügg et al. (2022); Rückin et al. (2023), F1-score (Rückin et al., 2023, 2024), or average precision (Chaplot et al., 2021), depending on the prediction task.
Due to the sequential nature of the AIPP problem, most works consider a well-performing AIPP approach not only to show strong evaluation metrics after a mission is over but also to ensure fast improvement during a mission as the remaining mission budget $B$ decreases. Trends in performance over mission time allow us to judge the efficiency of budget allocation, which is critical in AIPP as reflected by the constraints in Eq. 1. Common evaluation strategies include analyzing how the performance-based evaluation metrics evolve with the remaining mission budget, e.g., trajectory steps (Ott et al., 2023; Westheider et al., 2023; Viseras and Garcia, 2019; Cao et al., 2023a; Denniston et al., 2023a) or energy-based budgets (Popović et al., 2020; Rückin et al., 2022a; Hollinger and Sukhatme, 2014; Choudhury et al., 2020).
3.3 Benchmarks
Despite research efforts towards more generally applicable learning-based AIPP, most current methods are evaluated in application-dependent simulators and compared to task-dependent baseline algorithms. This results in a lack of standardized benchmark scenarios within the AIPP community. Benchmark scenarios are typically defined by two factors: (i) the choice of simulation environments and (ii) the choice of AIPP baseline algorithms against which the proposed methods are compared. Simulators fulfill a role analogous to datasets in the computer vision community as they define the problem setup in terms of the mission objective, evaluation environment, and robot configuration. AIPP baseline algorithms can be seen as paralleling widely established computer vision models, which are typically used to benchmark new models according to evaluation metrics.
Simulators for AIPP evaluation make various interdependent design choices leading to highly specific simulation environments. First, AIPP methods are typically evaluated on a single fixed task, e.g., the RockSample task (Ott et al., 2023; Choudhury et al., 2020), monitoring missions Popović et al. (2020); Rückin et al. (2022a); Westheider et al. (2023); Cao et al. (2023a); Choi and Cielniak (2021); Hollinger and Sukhatme (2014), or reconstruction (Zeng et al., 2022; Maboudi et al., 2023; Mendoza et al., 2020; Gazani et al., 2023; Dhami et al., 2023; Ran et al., 2022; Zhan et al., 2023; Yan et al., 2023). Second, simulators differ in the robot configuration defined by the robot type, e.g., autonomous underwater vehicle (AUV(Duecker et al., 2021), unmanned aerial vehicle (UAV(Popović et al., 2020), unmanned ground vehicle (UGV(Wei and Zheng, 2020), or a robot arm (Zeng et al., 2022), as well as the sensors, e.g., depth sensors (Cao et al., 2023b; Liu et al., 2023; Niroui et al., 2019; Hepp et al., 2018), thermal camera (Rückin et al., 2022a), point-based sensors Cao et al. (2023a); Hitz et al. (2017); Yang et al. (2023a); Marchant et al. (2014), or abstractions of sensor models interpreting raw sensor data (Popović et al., 2020; Chaplot et al., 2021; Denniston et al., 2023a; Georgakis et al., 2022b; Denniston et al., 2023b). Third, works vary in their assumptions about the structure of environments, e.g., 2D terrains (Hitz et al., 2017; Popović et al., 2020; Cao et al., 2023a; Rückin et al., 2022a; Ott et al., 2023; Choudhury et al., 2020) or 3D volumes (Cao et al., 2023b; Zeng et al., 2022; Rayas Fernández et al., 2022), and robot workspaces, e.g., obstacle-free 2D (Cao et al., 2023a; Velasco et al., 2020), obstacle-free 3D (Popović et al., 2020; Westheider et al., 2023; Rückin et al., 2022a; Denniston et al., 2023b; Rayas Fernández et al., 2022), or obstacle-aware 3D (Cao et al., 2023b; Zeng et al., 2022; Hepp et al., 2018) workspaces. Last, simulators vary according to the datasets they use, e.g., synthetic environments or randomly generated distributions (Popović et al., 2020; Cao et al., 2023a; Hitz et al., 2017; Rückin et al., 2022a; Westheider et al., 2023; Schmid et al., 2022; Bai et al., 2017; Saroya et al., 2020), more realistic simulators (Zeng et al., 2022; Rückin et al., 2023; Zacchini et al., 2023; Zhang et al., 2023), or real-world remote sensing datasets Popović et al. (2020); Rückin et al. (2023); Westheider et al. (2023); Rayas Fernández et al. (2022); Hitz et al. (2017); Hollinger and Sukhatme (2014); Caley et al. (2019). Sec. 4 and Sec. 5 provide a comprehensive overview of these aspects for each paper studied in our survey. In general, diverse non-unified simulator choices result in task- and application-dependent evaluation protocols. As a result, it is difficult to compare the performance of different AIPP algorithms in a generalizable and fair fashion under consistent conditions.
Despite the lack of standardized AIPP algorithms used for benchmarking new methods, some simple baseline algorithms have been established and used across different tasks and applications. These algorithms aim to foster exploration of the environment and use intuitive hand-crafted heuristics to optimize the evaluation metrics. We categorize approaches for exploration into two groups:
  1. 1.
    Geometric approaches, which pre-compute paths to maximize coverage of the initially unknown environment for uniform data collection. Examples include lawnmower-like grid patterns (Blum et al., 2019; Westheider et al., 2023; Galceran and Carreras, 2013) or spiral- or circle-like patterns Zeng et al. (2022); Popović et al. (2020); Zhang et al. (2023).
  2. 2.
    Random walk-like methods, which sample paths at random often combined with heuristics to efficiently manage the mission budget (Zurbrügg et al., 2022; Dai et al., 2022; Kumar et al., 2022; Sukkar et al., 2019; Liu et al., 2023).
In addition to these simple non-adaptive benchmarks, many works consider variants of well-established planning algorithms in the AIPP context. Classical planners used for evaluation include branch-and-bound techniques (Binney and Sukhatme, 2012) and sampling-based methods, e.g., Monte Carlo Tree Search (MCTSChoudhury et al. (2020), rapidly exploring information gathering (RIG) trees Hollinger and Sukhatme (2014), or rapidly exploring random trees (Karaman and Frazzoli, 2011). Other benchmarks include geometric strategies, e.g., greedy frontier-based exploration (Yamauchi, 1997), or optimization-based routines, e.g., Bayesian optimization (Gelbart et al., 2014) or the covariance matrix adaptation evolution strategy (CMA-ES(Hansen and Ostermeier, 2001). To enable adaptive replanning during a mission as new sensor measurements arrive, these algorithms are commonly implemented in a fixed- or receding-horizon fashion.
4 Learning Approaches
This section surveys the AIPP literature from a learning-based perspective. We provide a taxonomy classifying relevant works based on the underlying learning algorithm they use, discuss pertinent aspects in each category, and integrate them in our unified mathematical framework for AIPP.
4.1 Supervised Learning
Traditionally, the components of the general AIPP framework in Fig. 1 are realized using hand-crafted models or heuristics, e.g., forward sensor models (Popović et al., 2020; Lim et al., 2016; Karaman and Frazzoli, 2011; Rückin et al., 2022a) or hand-crafted information criteria (Hollinger and Sukhatme, 2014; Binney and Sukhatme, 2012; Best et al., 2019; Isler et al., 2016). This makes it difficult to incorporate prior knowledge and the expected distributions of geometry and information in complex scenarios, as well as to achieve adaptability in new environments. To address these drawbacks, various methods have been proposed to learn different elements of mapping and planning through supervised training from labeled data.
Given a dataset of $N$ training examples of the form $\{(x_{1},y_{1}),\ldots,(x_{N},y_{N})\}$, a supervised learning algorithm seeks to learn a ground truth function $g:X\rightarrow Y$, where $X$ and $Y$ are the input and output spaces, respectively, and $(x_{i},y_{i})\in X\times Y$ are dataset pairs. Table 3 provides a taxonomy of supervised learning applications in AIPP. These methods find diverse usage in both mapping and planning.
Supervised learning has been used for modeling environmental variables through Gaussian processes (Rayas Fernández et al., 2022; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Cao et al., 2023c; Marchant et al., 2014; Hitz et al., 2017; Popović et al., 2020; Hollinger and Sukhatme, 2014; Vivaldini et al., 2019; Viseras and Garcia, 2019; Yanes Luis et al., 2023; Choi and Cielniak, 2021; Wei and Zheng, 2020; Song et al., 2015; Denniston et al., 2023b). As described in Sec. 3.1, Gaussian processes are often used in Eq. 2 due to their probabilistic and non-parametric nature, making them suitable for uncertainty-based information gathering with spatially complex underlying feature mappings $F$. In this context, supervised learning is used to compute the ground truth function $g=F$ from a training dataset $(x_{i},z_{i})\in\xi\times\mathcal{F}$, with $X=\xi$ and $Y=\mathcal{F}$, obtained by taking stochastic measurements $z_{i}$ in the environment. Note that we use $x_{i}$ to denote general training inputs and $\mathbf{x}_{i}$ to denote the robot pose. In Gaussian process mapping, each $x_{i}$ corresponds to the pose $\mathbf{x}_{i}\in\xi$ introduced in Sec. 2. Example applications include mapping distributions of temperature (Popović et al., 2020; Westheider et al., 2023; Rückin et al., 2022a), light (Denniston et al., 2023a; Cao et al., 2023a), water properties (Hitz et al., 2017; Yanes Luis et al., 2023), and vegetation (Popović et al., 2020; Vivaldini et al., 2019). The learning process in implicit neural representations can be interpreted similarly. These methods learn a function that maps 3D coordinates $X=\xi$ to radiance values $Y=\mathcal{F}$, which represent color and light intensity, from a dataset of posed RGB (Jin et al., 2023; Pan et al., 2022; Zhan et al., 2023; Sünderhauf et al., 2023; Pan et al., 2023b) or RGB-D (Ran et al., 2022; Yan et al., 2023; Lee et al., 2022; Zhang et al., 2023) images.
For robotic exploration, several works focus on learning map completion from a partially observed map state, represented by its geometry (Shrestha et al., 2019; Zwecher et al., 2022; Georgakis et al., 2022a; Tao et al., 2023; Dhami et al., 2023; Li et al., 2023), topology (Saroya et al., 2020), or semantics (Schmid et al., 2023). At mission time step $t$, these methods predict the evolution of a probabilistic map representation $k$ steps into the future $\tilde{F}_{t+k}$. The dataset pairs $(x_{i},y_{i})$ used for learning map completion are application-dependent; for example, learning inputs can range from local observations (Schmid et al., 2023) to the current global map state (Li et al., 2020). The key advantage of such approaches is their ability to extrapolate incomplete maps by minimizing a specific loss during training. This extrapolation enables better reasoning with limited information and occlusions compared to classical heuristic strategies that plan solely based on the map state at a given time step.
Several works have also investigate the benefits of using supervised learning in the planning algorithm. One straightforward approach is to learn the information criterion $I(\cdot)$ in Eq. 1 from a dataset of actions and their associated information gains observed in previous data gathering missions $(x_{i},y_{i})=(a_{i},I(a_{i}))$, where $X=\mathcal{A}$ and $Y=\mathbb{R}$ (Hepp et al., 2018; Bai et al., 2017; Ly and Tsai, 2019). To solve the AIPP problem in Eq. 1 directly, the learning process can depend on pairs that connect map states to corresponding actions. Each action $a_{i}$ is created using a planning algorithm with full access to the ground truth data represented by $g$. Schmid et al. (2022) and Zacchini et al. (2023) propose learning a probabilistic model of informative views in the context of sampling-based planning. Alternatively, the next best viewpoint can be learned by reasoning about the shape of an object to be reconstructed (Vasquez-Gomez et al., 2021; Mendoza et al., 2020) or localization performance in the context of active SLAM (Hanlon et al., 2023). When the dataset pairs $(x_{i},y_{i})$ are generated through expert demonstrations, supervised learning can be thought of as a simple form of imitation learning (IL) as described in Sec. 4.3. By learning the information criterion $I(\cdot)$ from data, these approaches all bypass formulating it explicitly for online replanning.
A key requirement for all methods in Table 3 is the availability of reliable training data for model supervision. Some supervised learning methods in AIPP rely on open-source datasets, such as the indoor Matterport3D dataset (Georgakis et al., 2022a; Tao et al., 2023; Li et al., 2023) or outdoor 3D Street View dataset (Hepp et al., 2018). Works in environmental monitoring may use data from previously executed missions (Rayas Fernández et al., 2022; Hitz et al., 2017; Hollinger and Sukhatme, 2014; Caley et al., 2019), while others create custom synthetic environments approximating real-world scenarios (Schmid et al., 2022; Bai et al., 2017; Saroya et al., 2020; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Hitz et al., 2017; Popović et al., 2020, 2020; Viseras and Garcia, 2019; Zwecher et al., 2022; Song et al., 2015; Marchant et al., 2014). However, the lack of realistic labeled training data for AIPP problems remains an open issue restricting the generalizability of existing methods to new environments and domains. Furthermore, since supervised learning models are trained only on static data, their applicability is limited in dynamic environments with changing obstacles or terrains or the target information distribution varies over time. Finally, as most deep learning models do not consider uncertainty in the learning predictions, they lack natural mechanisms to handle sensor noise or other sources of uncertainty. In Sec. 4.4, we discuss how active learning (AL) techniques can mitigate some of these issues.
Algorithm keyword References
Architecture/method
Conditional variational autoencoder
(Schmid et al., 2022)
Convolutional neural network
(Schmid et al., 2022; Bai et al., 2017; Ly and Tsai, 2019; Caley and Hollinger, 2020; Caley et al., 2019; Zwecher et al., 2022; Hepp et al., 2018; Vasquez-Gomez et al., 2021; Mendoza et al., 2020; Li et al., 2023; Saroya et al., 2020; Georgakis et al., 2022a; Ramakrishnan et al., 2020; Li et al., 2023; Pan et al., 2023b, a)
Gaussian process
(Rayas Fernández et al., 2022; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Cao et al., 2023c; Marchant et al., 2014; Hitz et al., 2017; Popović et al., 2020; Hollinger and Sukhatme, 2014; Vivaldini et al., 2019; Viseras and Garcia, 2019; Yanes Luis et al., 2023; Choi and Cielniak, 2021; Wei and Zheng, 2020; Song et al., 2015; Denniston et al., 2023b)
Implicit neural representation
(Jin et al., 2023; Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023; Sünderhauf et al., 2023; Yan et al., 2023; Lee et al., 2022; Zhang et al., 2023; Pan et al., 2023b)
Kernel density estimation
(Zacchini et al., 2023)
Occupancy prediction network
(Tao et al., 2023)
Scene completion network
(Schmid et al., 2023)
Transformer
(Dhami et al., 2023; Hanlon et al., 2023)
Variational autoencoder
(Shrestha et al., 2019)
Training objective
3D semantic scene completion
(Schmid et al., 2023)
2D occupancy map completion
(Shrestha et al., 2019; Zwecher et al., 2022; Georgakis et al., 2022a; Ramakrishnan et al., 2020)
3D occupancy map completion
(Tao et al., 2023)
Environment topology
(Saroya et al., 2020)
Environmental variable
(Rayas Fernández et al., 2022; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Cao et al., 2023c; Marchant et al., 2014; Hitz et al., 2017; Popović et al., 2020; Hollinger and Sukhatme, 2014; Vivaldini et al., 2019; Popović et al., 2020; Viseras and Garcia, 2019; Yanes Luis et al., 2023; Choi and Cielniak, 2021; Caley and Hollinger, 2020; Wei and Zheng, 2020; Song et al., 2015; Denniston et al., 2023b)
Grasp affordance
(Zhang et al., 2023)
Informative view distribution
(Schmid et al., 2022; Zacchini et al., 2023)
Localization score
(Hanlon et al., 2023)
Minimum view subset for coverage
(Pan et al., 2023a)
Next-best view
(Vasquez-Gomez et al., 2021; Mendoza et al., 2020)
Point cloud shape completion
(Dhami et al., 2023)
Radiance field
(Jin et al., 2023; Pan et al., 2022; Zhang et al., 2023; Ran et al., 2022; Zhan et al., 2023; Lee et al., 2022; Sünderhauf et al., 2023; Yan et al., 2023)
Required number of views
(Pan et al., 2023b)
Target point of interest
(Caley et al., 2019)
Unexplored navigable area beyond map frontiers
(Li et al., 2023)
Utility/reward
(Hepp et al., 2018; Schmid et al., 2022; Bai et al., 2017; Ly and Tsai, 2019)
Mission objective
Coverage
(Schmid et al., 2023, 2022; Saroya et al., 2020; Ly and Tsai, 2019; Vasquez-Gomez et al., 2021; Mendoza et al., 2020; Zwecher et al., 2022; Tao et al., 2023; Li et al., 2023)
Estimate quantile values
(Rayas Fernández et al., 2022)
Find maximum field value
(Song et al., 2015; Denniston et al., 2023b)
Maximize grasp quality
(Zhang et al., 2023)
Maximize localization accuracy
(Hanlon et al., 2023)
Maximize probability of improvement/expected improvement
(Denniston et al., 2023b)
Maximize observed surfaces
(Dhami et al., 2023; Pan et al., 2023a)
Minimize explored area to target point of interest
(Caley et al., 2019)
Minimize map error
(Viseras and Garcia, 2019; Choi and Cielniak, 2021; Caley and Hollinger, 2020; Ramakrishnan et al., 2020)
Minimize map uncertainty
(Hepp et al., 2018; Bai et al., 2017; Shrestha et al., 2019; Zacchini et al., 2023; Hollinger and Sukhatme, 2014; Yanes Luis et al., 2023; Georgakis et al., 2022a; Wei and Zheng, 2020; Denniston et al., 2023b)
Minimize map uncertainty and robot localization uncertainty
(Popović et al., 2020)
Minimize map uncertainty in high-interest areas
(Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Marchant et al., 2014; Hitz et al., 2017; Popović et al., 2020; Hollinger and Sukhatme, 2014)
Minimize map uncertainty in unclassified areas
(Vivaldini et al., 2019)
Minimize model uncertainty
(Jin et al., 2023; Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023; Lee et al., 2022; Sünderhauf et al., 2023; Yan et al., 2023)
Minimize target localization uncertainty
(Cao et al., 2023c)
Object reconstruction based on PSNR
(Pan et al., 2023b)
Point-goal navigation
(Georgakis et al., 2022a; Ramakrishnan et al., 2020)
Training data
3D Street View dataset
(Hepp et al., 2018)
Custom CAD models
(Vasquez-Gomez et al., 2021; Mendoza et al., 2020; Ran et al., 2022; Yan et al., 2023; Pan et al., 2023a)
Custom synthetic dataset
(Schmid et al., 2022; Bai et al., 2017; Saroya et al., 2020; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Hitz et al., 2017; Popović et al., 2020, 2020; Viseras and Garcia, 2019; Zwecher et al., 2022; Song et al., 2015; Marchant et al., 2014)
DTU dataset
(Jin et al., 2023)
Heuristic parameter setting
(Cao et al., 2023c; Yanes Luis et al., 2023)
Gibson dataset
(Ramakrishnan et al., 2020)
INRIA Aerial Image Labeling dataset
(Ly and Tsai, 2019)
KTH dataset
(Shrestha et al., 2019; Caley et al., 2019)
LLFF dataset
(Pan et al., 2022; Sünderhauf et al., 2023)
Matterport3D dataset
(Georgakis et al., 2022a; Tao et al., 2023; Li et al., 2023; Hanlon et al., 2023)
NeRF/Synthetic-NeRF dataset
(Pan et al., 2022; Ran et al., 2022; Lee et al., 2022; Sünderhauf et al., 2023; Zhan et al., 2023)
NYU dataset
(Schmid et al., 2023)
Real-world environmental data
(Rayas Fernández et al., 2022; Hitz et al., 2017; Hollinger and Sukhatme, 2014; Caley et al., 2019; Wei and Zheng, 2020; Denniston et al., 2023b; Vivaldini et al., 2019)
Realistic simulator
(Zacchini et al., 2023; Zhang et al., 2023)
Regional Ocean Modeling System (ROMS) dataset
(Caley and Hollinger, 2020)
ShapeNet dataset
(Dhami et al., 2023; Jin et al., 2023; Pan et al., 2023b)
Tanks&Temples dataset
(Zhan et al., 2023)
Table 3: Supervised learning methods for AIPP.
4.2 Reinforcement Learning
[…]
RL is a branch of machine learning that equips an agent to learn from an environment by interacting with it and receiving feedback in terms of rewards or penalties. In AIPP, reinforcement learning (RL) has emerged as an effective learning methodology, enabling a system to adapt its behavior based on an evolving understanding of its environment.
The problem of AIPP can be framed as a partially observable Markov decision process (POMDP)  (Kochenderfer et al., 2022). Formally, a POMDP is defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{O},T,Z,R\rangle$, where:
  • 1.
    $\mathcal{S}$ is the state space;
  • 2.
    $\mathcal{A}$ is the action space;
  • 3.
    $\mathcal{O}$ is the observation space;
  • 4.
    $T(s^{\prime}\mid s,a)$ is the state transition function;
  • 5.
    $Z(o\mid s^{\prime},a)$ is the observation function;
  • 6.
    $R$ is the immediate reward function.
In many real-world scenarios, the agent cannot directly observe the true state of the environment. Instead, it receives observations that provide partial information about the true state. In the context of AIPP, the true state of the environment is the actual physical attributes of the environment captured by $F(\mathbf{x})$ in Sec. 2, e.g., the locations of obstacles and areas of interest, while the observation $z$ is the robot’s current measurement of these attributes based on sensor readings.
The state space $\mathcal{S}$ is defined as the set of all possible feature mappings $F$ in the environment and agent-specific information, e.g., its position or battery level. The action space $\mathcal{A}$ consists of all possible actions $a_{i}\in\mathcal{A}$ that the robot could take, e.g., moving to a given location, using a specific sensor, or returning to a charging station. The observation space $\mathcal{O}$ corresponds to the set of all possible sensor readings and derived interpretations the robot could have about the environment, e.g., semantic segmentation of RGB images.
The state transition function $T(s^{\prime}\,|\,s,a)$ describes the probabilistic dynamics of the robot’s state in the context of the AIPP problem. When the robot executes an action $a\in\mathcal{A}$ in its current state $s\in\mathcal{S}$, it moves to a new state $s^{\prime}\in\mathcal{S}$. The inherent uncertainty of action outcomes and their impact on the environment is captured by the transition function. This function is used in the belief update, which refines the robot’s understanding of the environment feature mapping $F$. Additionally, the observation function $Z(o\,|\,s^{\prime},a)$ provides the likelihood of perceiving an observation $o\in\mathcal{O}$ subsequent to action $a$ and arriving at state $s^{\prime}$. This observation model, which is probabilistically linked to the feature mapping $F$ through $z_{t}\sim p(z\,|\,F)$, is used to update the robot’s belief state using Eq. 2. After executing action $a$ and observing $o$, the belief over the feature mapping $F$ is updated. This update enables the robot to better predict environmental features, which in turn aids in refining its path planning.
The reward function $R(s,a)$ captures the trade-off between maximizing information gain, as described by the criterion $I(\cdot)$, and adhering to a budget constraint $C(\psi)\leq B$. While the objective function $I(\cdot)$ evaluates a sequence of actions, the reward function $R(s,a)$ evaluates individual actions. That is, the information criterion $I(\cdot)$ in AIPP accumulates the rewards $R(s,a)$ over a sequence of actions, and can additionally incorporate a discount factor $\gamma$ to reflect that future rewards are less valuable than immediate rewards:
$I(\psi)=\sum_{t=1}^{N}\gamma^{t-1}R(s_{t},a_{t})\,.$ (3)
In AIPP, the discount factor is typically set to $\gamma=1$ for finite horizons with budget constraints.
The policy is a function $\pi:\mathcal{S}\rightarrow\mathcal{A}$. The concept of a policy and the general action sequence $\psi$ in Eq. 1 are linked by $\psi=(\pi(s_{1}),\dots,\pi(s_{N}))$, where the states evolve based on the previous state and action according to $T(s^{\prime}\,|\,s,a)$. The optimal policy can be queried to select an action sequence $\psi^{*}$ maximizing the total expected discounted return, and hence the information gain, while being within the allowed budget.
Any POMDP can be viewed as an MDP that uses beliefs as states, also called a belief-state MDP or belief MDP. The state space of a belief MDP is the set of all beliefs $\mathcal{B}$. The action space is equivalent to that of the POMDP (Kochenderfer et al., 2022). To connect this directly to Sec. 3, in AIPP, the robot’s belief about the true state of the environment is denoted by $\tilde{F}$ in Sec. 3.1. This belief evolves as the robot gathers information about its environment, which starts as partially unknown and is gradually discovered through sensing and exploration. As discussed in Sec. 3.1, the structure of this belief involves the robot creating an environment map from the sensor measurements. The robot’s beliefs about the feature mappings, represented by $\tilde{F}$, are updated as it collects measurements and takes actions. These beliefs are then used to quantify the change in information criterion $I(\cdot)$. The performance of a specific policy can be evaluated using the metrics described in Sec. 3.2 based on the posterior map belief $\tilde{F}$.
A variety of RL-based approaches have been explored in the literature. In Table 4, we distinguish the methods primarily by how they formulate each component of the POMDP architecture as well as how they learn policies.
All of the methods surveyed include some aspect of the current robot information in the state. The robot pose is often used (Cao et al., 2023b; Lodel et al., 2022; Zeng et al., 2022; Marchant et al., 2014; Denniston et al., 2023b), while other methods include additional robot information, e.g., energy level (Choudhury et al., 2020; Ott et al., 2023), viewing direction (Chen et al., 2022), and operational status (Bayerlein et al., 2021). In addition to current robot information, all methods store some form of a spatial map in the state. The most popular spatial maps are Gaussian processes (Cao et al., 2023a, c; Yang et al., 2023a; Wei and Zheng, 2020; Rückin et al., 2022a; Ott et al., 2023; Yanes Luis et al., 2023; Viseras and Garcia, 2019; Choi and Cielniak, 2021; Marchant et al., 2014; Denniston et al., 2023b) and occupancy grids (Yang et al., 2023b; Niroui et al., 2019; Chen et al., 2020; Lodel et al., 2022; Chen et al., 2019; Westheider et al., 2023; Liu et al., 2023; Gao and Zhang, 2022; Zwecher et al., 2022; Chen et al., 2022, 2021; Cao et al., 2023b; Sukkar et al., 2019), both of which are described in Sec. 3.1. Another common component of the state is the previous position and sensor history. The history components can range from the executed trajectory (Cao et al., 2023a) to the previous graph history (Cao et al., 2023c) as well as the previous observations (Viseras and Garcia, 2019; Denniston et al., 2023a). Some methods implicitly capture the history of robot poses and measurements in the belief state action-observation history (Choudhury et al., 2020; Ott et al., 2023) or by capturing the previously explored areas in the occupancy grid (Chen et al., 2021; Westheider et al., 2023; Gao and Zhang, 2022).
[…]
Continuous action spaces often significantly increase the computational complexity of POMDPs. In many cases, discrete action spaces are simpler to work with. When the spatial map is represented as an abstract graph, a common choice is for movement commands to be represented as movement from one graph node to another via the connecting edge (Cao et al., 2023a, b, c; Yang et al., 2023a; Wei and Zheng, 2020; Rückin et al., 2022a; Ott et al., 2023; Choudhury et al., 2020; Westheider et al., 2023; Yanes Luis et al., 2023; Viseras and Garcia, 2019; Viseras et al., 2021; Bayerlein et al., 2021; Denniston et al., 2023a; Duecker et al., 2021; Denniston et al., 2023b; Gao and Zhang, 2022; Best et al., 2019; Zwecher et al., 2022). Another common approach relies on frontiers, which represent the boundary between known and unknown space in the environment and can be seen as a special case of a graph node action. The use of frontiers as movement actions does not necessarily require the spatial map to be represented as an abstract graph (Niroui et al., 2019; Chen et al., 2020, 2021; Liu et al., 2023). Other discrete actions include selecting from a discrete set of sensors to employ (Choudhury et al., 2020; Ott et al., 2023) as well as motion primitive commands (Marchant et al., 2014; Song et al., 2015) which can be seen as a discretization of continuous actions where the agent is given a set of potential trajectories to execute which correspond to actions.
The reward function is the most problem-dependent component of the formulation. One of the core challenges that arise in the POMDP framework and, consequently, in AIPP, is the exploration-exploitation trade-off. The agent must select actions that balance between exploring unknown areas of the environment to increase its knowledge (exploration) and exploiting its current knowledge to find areas of interest within the constraints of its available resources (exploitation). This trade-off is intrinsically linked to the AIPP objective of maximizing the information criterion while respecting the cost constraint. Hence, it is a critical factor driving the learning process of the RL agent. Essentially, all AIPP methods use some measure of informativeness as the objective as discussed in Sec. 3.2. Map uncertainty reduction in high-interest areas is a very common RL reward in the AIPP literature (Yang et al., 2023a; Ott et al., 2023; Rückin et al., 2022a; Westheider et al., 2023; Viseras et al., 2021; Denniston et al., 2023a; Marchant et al., 2014; Duecker et al., 2021; Denniston et al., 2023b; Song et al., 2015). Some approaches use only map uncertainty reduction to spread measurements across the environment as uniformly as possible and minimize the estimation error Cao et al. (2023a); Niroui et al. (2019); Cao et al. (2023c); Ott et al. (2023); Yanes Luis et al. (2023); Viseras and Garcia (2019); Choi and Cielniak (2021); Cao et al. (2023b). Coverage is a common surrogate metric for map uncertainty reduction since it only relies on computing the area uncovered as the robot travels through the environment, which is often faster than computing the Gaussian process variance online (Zeng et al., 2022; Liu et al., 2023; Kumar et al., 2022; Bayerlein et al., 2021; Sukkar et al., 2019; Best et al., 2019; Zwecher et al., 2022; Chen et al., 2022). While coverage can perform well on the map uncertainty reduction objective, it cannot explicitly reason about high-interest areas. The mutual information between observations and landmarks is another common way to measure map uncertainty reduction (Yang et al., 2023b; Lodel et al., 2022; Wei and Zheng, 2020; Chen et al., 2019). Some methods also explicitly incorporate distance-based metrics in the reward function (Chen et al., 2020, 2021; Cao et al., 2023b; Lodel et al., 2022; Gao and Zhang, 2022). For example, Chen et al. (2020, 2021) and Cao et al. (2023b) assign greater rewards for actions that achieve the same uncertainty reduction in less distance.
Optimally solving POMDPs is computationally challenging (Kaelbling et al., 1998; Kochenderfer et al., 2022). However, practical solutions can be obtained through various RL algorithms. We categorize AIPP policies based on their deployment inference strategies, distinguishing between zero-shot inference and expected discounted $m$-step search policies. Zero-shot inference policies are trained offline, considering the sequence of actions. The policy is trained to account for the cascading impact of each action on subsequent decisions, taking into consideration the consequences of actions and future dependencies during the offline training process. Upon deployment, these policies take the current belief state—updated to reflect the cumulative knowledge of the environment as per Eq. 2—and directly output the next best action without additional planning. Conversely, policies using an expected discounted $m$-step search are dynamic in the sense that they actively consider a sequence of potential future actions during deployment. Instead of reacting to the current belief state, this approach projects forward, evaluating a decision tree of potential action sequences up to $m$ steps ahead and factoring in the anticipated updates to the belief state for each action. The policy then chooses an action based on the strategy that maximizes the total expected reward over this lookahead horizon.
Description References
State
Current robot information
Yang et al. (2023b); Cao et al. (2023a); Niroui et al. (2019); Chen et al. (2020, 2021); Cao et al. (2023b); Lodel et al. (2022); Cao et al. (2023c); Yang et al. (2023a); Zeng et al. (2022); Wei and Zheng (2020); Rückin et al. (2022a); Chen et al. (2019); Ott et al. (2023); Choudhury et al. (2020); Westheider et al. (2023); Yanes Luis et al. (2023); Liu et al. (2023); Viseras and Garcia (2019); Viseras et al. (2021); Bartolomei et al. (2021); Kumar et al. (2022); Choi and Cielniak (2021); Bayerlein et al. (2021); Gazani et al. (2023); Denniston et al. (2023a); Marchant et al. (2014); Duecker et al. (2021); Denniston et al. (2023b); Gao and Zhang (2022); Sukkar et al. (2019); Best et al. (2019); Hüttenrauch et al. (2019); Zwecher et al. (2022); Chen et al. (2022); Qie et al. (2019); Fan et al. (2020); Arora et al. (2019); Bucher et al. (2021); Singh et al. (2009)
Observation history
Ye et al. (2018); Lodel et al. (2022); Cao et al. (2023c); Choudhury et al. (2020); Westheider et al. (2023); Liu et al. (2023); Viseras and Garcia (2019); Gazani et al. (2023); Denniston et al. (2023a); Bucher et al. (2021); Singh et al. (2009)
Robot history
Cao et al. (2023a); Chen et al. (2020, 2021); Cao et al. (2023c); Gazani et al. (2023); Denniston et al. (2023a); Gao and Zhang (2022)
Spatial map
Yang et al. (2023b); Cao et al. (2023a); Niroui et al. (2019); Chen et al. (2020, 2021); Cao et al. (2023b); Lodel et al. (2022); Cao et al. (2023c); Yang et al. (2023a); Zeng et al. (2022); Wei and Zheng (2020); Rückin et al. (2022a); Chen et al. (2019); Ott et al. (2023); Choudhury et al. (2020); Westheider et al. (2023); Yanes Luis et al. (2023); Liu et al. (2023); Viseras and Garcia (2019); Viseras et al. (2021); Bartolomei et al. (2021); Kumar et al. (2022); Choi and Cielniak (2021); Bayerlein et al. (2021); Gazani et al. (2023); Denniston et al. (2023a); Marchant et al. (2014); Duecker et al. (2021); Denniston et al. (2023b); Gao and Zhang (2022); Sukkar et al. (2019); Best et al. (2019); Hüttenrauch et al. (2019); Zwecher et al. (2022); Chen et al. (2022); Qie et al. (2019); Fan et al. (2020); Arora et al. (2019); Singh et al. (2009)
Map
2D Gaussian process
Cao et al. (2023a, c); Yang et al. (2023a); Wei and Zheng (2020); Rückin et al. (2022a); Ott et al. (2023); Yanes Luis et al. (2023); Viseras and Garcia (2019); Choi and Cielniak (2021); Marchant et al. (2014); Denniston et al. (2023b)
2D occupancy grid
Yang et al. (2023b); Niroui et al. (2019); Chen et al. (2020); Lodel et al. (2022); Chen et al. (2019); Westheider et al. (2023); Liu et al. (2023); Gao and Zhang (2022); Zwecher et al. (2022); Chen et al. (2022)
3D occupancy grid
Chen et al. (2021); Cao et al. (2023b); Sukkar et al. (2019)
3D semantic occupancy map
(Zeng et al., 2022; Bartolomei et al., 2021)
Bayesian network
Arora et al. (2019)
Factor graph
Denniston et al. (2023a)
Gaussian Markov random field
Duecker et al. (2021)
Landmark locations and uncertainties
Chen et al. (2020, 2021); Bartolomei et al. (2021); Fan et al. (2020)
Location graph
Best et al. (2019); Choudhury et al. (2020); Hüttenrauch et al. (2019); Qie et al. (2019); Singh et al. (2009)
Action
Continuous action
Yang et al. (2023b); Lodel et al. (2022); Bartolomei et al. (2021); Sukkar et al. (2019); Hüttenrauch et al. (2019); Fan et al. (2020); Denniston et al. (2023b); Bucher et al. (2021)
Discrete action
Zeng et al. (2022); Chen et al. (2019); Ott et al. (2023); Viseras et al. (2021); Kumar et al. (2022); Choi and Cielniak (2021); Bayerlein et al. (2021); Gazani et al. (2023); Marchant et al. (2014); Chen et al. (2022); Qie et al. (2019); Arora et al. (2019); Ye et al. (2018)
Frontier
Niroui et al. (2019); Chen et al. (2020, 2021); Liu et al. (2023)
Graph node
Cao et al. (2023a, b, c); Yang et al. (2023a); Wei and Zheng (2020); Rückin et al. (2022a); Ott et al. (2023); Choudhury et al. (2020); Westheider et al. (2023); Yanes Luis et al. (2023); Viseras and Garcia (2019); Viseras et al. (2021); Bayerlein et al. (2021); Denniston et al. (2023a); Duecker et al. (2021); Denniston et al. (2023b); Gao and Zhang (2022); Best et al. (2019); Zwecher et al. (2022); Singh et al. (2009)
Policy
Expected discounted $m$-step reward
Rückin et al. (2022a); Ott et al. (2023); Choudhury et al. (2020); Denniston et al. (2023a); Marchant et al. (2014); Duecker et al. (2021); Denniston et al. (2023b); Sukkar et al. (2019); Best et al. (2019); Arora et al. (2019); Bucher et al. (2021)
Zero-shot inference
Yang et al. (2023b); Cao et al. (2023a); Niroui et al. (2019); Chen et al. (2020, 2021); Cao et al. (2023b); Lodel et al. (2022); Cao et al. (2023c); Yang et al. (2023a); Zeng et al. (2022); Wei and Zheng (2020); Chen et al. (2019); Westheider et al. (2023); Yanes Luis et al. (2023); Liu et al. (2023); Viseras and Garcia (2019); Viseras et al. (2021); Bartolomei et al. (2021); Kumar et al. (2022); Choi and Cielniak (2021); Bayerlein et al. (2021); Gazani et al. (2023); Gao and Zhang (2022); Hüttenrauch et al. (2019); Zwecher et al. (2022); Chen et al. (2022); Qie et al. (2019); Fan et al. (2020); Denniston et al. (2023b); Ye et al. (2018); Singh et al. (2009)
Reward
Belief state mode
Choudhury et al. (2020)
Coverage
Zeng et al. (2022); Liu et al. (2023); Kumar et al. (2022); Bayerlein et al. (2021); Sukkar et al. (2019); Best et al. (2019); Zwecher et al. (2022); Chen et al. (2022); Bucher et al. (2021)
Maximize inter-agent distance
(Hüttenrauch et al., 2019)
Maximize mutual information of observations and landmarks
Yang et al. (2023b); Lodel et al. (2022); Wei and Zheng (2020); Chen et al. (2019)
Minimize landmark localization uncertainty
Chen et al. (2020, 2021); Bartolomei et al. (2021); Gazani et al. (2023)
Minimize map uncertainty
Cao et al. (2023a); Niroui et al. (2019); Cao et al. (2023c); Ott et al. (2023); Yanes Luis et al. (2023); Viseras and Garcia (2019); Choi and Cielniak (2021); Cao et al. (2023b)
Minimize map uncertainty in high-interest areas
Yang et al. (2023a); Ott et al. (2023); Rückin et al. (2022a); Westheider et al. (2023); Viseras et al. (2021); Denniston et al. (2023a); Marchant et al. (2014); Duecker et al. (2021); Denniston et al. (2023b); Arora et al. (2019); Singh et al. (2009)
Size of bounding box for target detection
Ye et al. (2018)
Travel distance
Chen et al. (2020, 2021); Cao et al. (2023b); Lodel et al. (2022); Gao and Zhang (2022); Qie et al. (2019)
Travel time
Fan et al. (2020)
Table 4: Reinforcement learning methods for AIPP.
Yang et al. (2023b) use an attention-based multi-layer perceptron (MLP) which is trained using proximal policy optimization (PPO) to produce a zero-shot inference action of continuous longitudinal and lateral velocity with the goal of maximizing the mutual information between observations and landmark states. Cao et al. (2023a) take a slightly different approach by using an attention-based graph neural network and long short-term memory (LSTM) network with PPO to select the next graph node. Rückin et al. (2022a) use AlphaZero by combining Monte Carlo tree search (MCTS) with a convolutional neural network (CNN) to learn information-rich actions in adaptive data gathering missions. Chen et al. (2019) uses deep Q-network (DQN) with a CNN to select the next location to visit from a set of randomly sampled 2D positions with the goal of maximizing the mutual information between received observations and the occupancy grid map. Chen et al. (2021) uses the advantage actor-critic (A2C) algorithm with a graph neural network to select the next best frontier to visit with the goal of maximizing landmark uncertainty reduction and minimizing travel distance. Cao et al. (2023b) uses the soft actor-critic (SAC) algorithm with an attention-based graph neural network to output the next graph node to visit with the goal of maximizing the number of observed frontiers while minimizing travel distance. Other methods mainly differ in the training algorithms and the network architectures used with the most popular being PPO, A2C, SAC, and CNN, DQN, LSTM, respectively (Yang et al., 2023b; Cao et al., 2023a; Niroui et al., 2019; Chen et al., 2020, 2021; Cao et al., 2023b; Lodel et al., 2022; Cao et al., 2023c; Chen et al., 2019; Liu et al., 2023; Gazani et al., 2023; Zwecher et al., 2022).
In summary, RL offers a powerful framework for AIPP. By learning to optimize actions for maximum information gain, RL-based systems can adaptively and effectively explore unknown environments. However, RL algorithms also face challenges such as data-intensive training, sensitivity to reward design, and computational demands, which can affect their efficiency and generalizability in real-world robotics scenarios. Existing RL-based AIPP literature addresses these issues through advanced neural network architectures, sophisticated training algorithms, careful reward and policy design, and robust map representations. Additionally, most of these approaches are tested on environments very similar to those used for training. A promising direction for future work is thus to develop methods that can generalize to new environments.
4.3 Imitation Learning
[…]
A simple form of imitation learning (IL) can be thought of as a supervised learning problem from Sec. 4.1 typically referred to as behavioral cloning Pomerleau (1991). The objective is to train a stochastic policy $\pi:\mathcal{S}\to\mathcal{A}$ where the policy $\pi_{\mathbf{\theta}}$ is parameterized by $\mathbf{\theta}$. The training objective is to maximize the likelihood of actions from a dataset $\mathcal{D}$ of expert state-action pairs:
[…]
This approach, while straightforward, is limited by the quality and quantity of the expert demonstrations available. Moreover, it often generalizes poorly to states not encountered in the training dataset, leading to suboptimal or even erroneous behaviors under novel circumstances. More sophisticated IL strategies have been developed to address these limitations. These include methods combining IL with RL to refine policies beyond the initial demonstrations, approaches that incorporate active learning (AL) to query experts for demonstrations in states where the learner is uncertain, and algorithms that utilize inverse RL to infer the underlying reward function implied by the expert’s behavior.
It is important to note the similarities and differences between IL and RL. Both IL and RL seek to find an optimal policy for a given task. However, the key distinction lies in how they learn this policy. RL aims to learn the policy by maximizing the expected discounted return, as illustrated in Eq. 3. In contrast, IL, as depicted in Eq. 4, assumes access to samples from an expert policy and focuses on mimicking this behavior. This approach treats the expert policy as a black box, learning to replicate its decisions without necessarily understanding the underlying motivations. In IL, the reliance on expert demonstrations aligns closely with the concept of an active learning (AL) oracle discussed in Sec. 4.4, which provides guidance in the form of expert input, albeit in a more passive manner.
IL has been effectively applied in robotic scenarios, including indoor autonomous exploration, path planning, and multi-robot coordination. Liu et al. (2023) introduce the Learning to Explore (L2E) framework, combining IL with deep RL for UGVs in indoor environments. This approach pretrains exploration policies using IL and then refines them with RL, leading to improved sample efficiency and training speed and outperforming heuristic methods in diverse environments.
Choudhury et al. (2018) propose a data-driven IL framework to train planning policies by imitating a clairvoyant oracle with ground truth knowledge of the world map. This approach efficiently trains policies based on partial information for tasks like AIPP and motion planning, showing significant performance gains over state-of-the-art algorithms, even in real UAV applications. Reinhart et al. (2020) develop a learning-based path planning approach for UAVs in unknown subterranean environments. They use a graph-based path planner as a training expert for IL, resulting in a policy that guides autonomous exploration with reduced computational costs and less reliance on consistent, online reconstructed maps.
Li et al. (2020) address decentralized multi-robot path planning by proposing a combined model that synthesizes local communication and decision-making policies. Their architecture, comprising a convolutional neural network and a graph neural network, was trained to imitate an expert. This model demonstrates its effectiveness in decentralized planning with local communication and observations in cluttered workspaces. Tzes et al. (2023) tackles multi-robot AIPP with their Information-aware Graph Block Network (I-GBNet). Trained via IL, this network aggregates information over a communication graph and offers sequential decision-making in a distributed manner. Its scalability and robustness were validated in large-scale experiments involving dynamic target tracking and localization. Lastly, Dai et al. (2022) explores indoor environments using an RGB-D camera and generative adversarial imitation learning (GAIL). They modified the ORB-SLAM2 method for enhanced navigation and trained a camera view planning policy with GAIL, yielding efficient exploration and reducing tracking failure in indoor environments.
These studies show the versatility and effectiveness of IL in a wide range of robotic applications, demonstrating its potential to enhance efficiency, decision-making, and adaptability in varied and challenging environments. However, significant challenges and areas for future research remain. Key challenges include improving the generalizability of learned behaviors to unseen environments, enhancing the robustness of IL algorithms against dynamic and unpredictable real-world conditions, and developing more efficient methods for collecting and using expert demonstrations. Further work is also needed to integrate IL with other learning paradigms, such as RL, to address complex multi-agent systems and high-dimensional state spaces.
4.4 Active Learning
The problem of AIPP can also be seen as an active learning (AL) process, whereby the learner uses information measures to choose actions for collecting useful or descriptive data (Taylor et al., 2021). Here, we mathematically connect the general concept of AL to robotic AIPP theory. We then draw parallels between different components of AL and the learning-based AIPP methods presented in the previous subsections. Lastly, we elaborate on an emerging line of work in AIPP that uses AL objectives to improve sensor models in new environments.
AL addresses the task of finding the most informative data in a set of unlabeled samples such that a model is maximally improved if that sample is labeled and added to the existing training dataset (Settles, 2009). We define the AL problem as follows:
  • 1.
    $\mathcal{Q}$ is the space of queries $q_{i}\in\mathcal{Q}$ and $N_{q}\in\mathbb{N}$ is the number of sampled queries;
  • 2.
    $\mathcal{L}$ is the space of targets $l_{i}\in\mathcal{L}$ corresponding to query $q_{i}$;
  • 3.
    $\theta:\mathcal{Q}\to\mathcal{L}$ is the model predicting a target $\theta(q_{i})=l_{i}$ given a query $q_{i}$;
  • 4.
    $P:\Theta\to\mathbb{R}$ is the performance function evaluating a model $\theta\in\Theta$;
  • 5.
    $O:\mathcal{Q}\to\mathcal{L}$ is an oracle that can be queried to generate a target $O(q_{i})=l_{i}$ for a specific query $q_{i}$;
  • 6.
    $q=(q_{1},\ldots,q_{i})\in\mathcal{Q}^{i}$, $l=(O(q_{1}),\ldots,O(q_{i}))\in\mathcal{L}^{i}$ are collected queries and targets up to iteration $i\leq N_{q}$;
  • 7.
    $A:\Theta\times\mathcal{Q}\to\mathbb{R}$ is the acquisition function estimating a query’s $q_{i}$ effect on the model performance $P(\theta)$ after re-training on collected training data pairs $q$ and $l$.
We aim to maximize the model’s performance with a low number of sampled queries answered by the oracle. Formally, AL solves the following optimization problem:
$q^{*}=\operatorname*{argmax}_{q\in\mathcal{Q}^{N_{q}}}\sum_{i=1}^{N_{q}}A(\theta,q_{i})\,.$ (5)
The AIPP problem in Eq. 1 and AL in Eq. 5 are closely related. In the AL setup, the costs $C(q)=\lvert q\rvert$ of executing a sequence of queries $q$ is given by the number of queries $\lvert q\rvert$ in $q$. The optimization constraints of Eq. 5 are implicitly determined by $C(q)\leq N_{q}$. The AIPP information criterion $I$ translates to $I(q)=\sum_{i=1}^{N_{q}}A(\theta,q_{i})$ by substituting the AL acquisition function as the mission objective.
In the classical environmental monitoring or exploration problem setup, the environment $\xi$ is characterized by the set of available queries $\mathcal{Q}$, e.g., robot poses $q_{i}=\mathbf{x}_{i}\in\xi$. The feature space $\mathcal{F}$ encompasses the set of targets $\mathcal{L}$ establishing a correspondence between feature mappings $F:\xi\to\mathcal{F}$ and the model $\theta$. Further, we cannot directly quantify the true effect of a sequence of queries $q$ on the model performance $P(\theta)$ as this requires re-training on not-yet-acquired oracle-generated targets. Instead, similar to the belief update of $\tilde{F}$ over the true feature mapping $F$ in Eq. 2, at each iteration $i$, AL methods leverage already collected queries $\{q_{1},\ldots,q_{i}\}$ and targets $\{O(q_{1}),\ldots,O(q_{i})\}$ to update the model $\theta$ and ensure maximally informed sequential query selection. We establish this theoretical connection between AIPP and the AL problem to show that the AIPP methods introduced in Sec. 4.1, 4.2 and 4.3 can be viewed as methods for AL in the context of embodied robotic applications. In the following, we identify the models, queries, oracles, and acquisition functions used in different learning-based approaches from an AL perspective.
The model could have any parametric or non-parametric form. As discussed in Sec. 4.1, common models for learning a spatial representation $F:\xi\to\mathcal{F}$ of a feature space $\mathcal{F}$ in an environment $\xi$ are non-parametric Gaussian processes or implicit neural representations to probabilistically model $O:\xi\to\mathcal{F}$. In this case, a query $q_{i}$ is defined by a specific location or area in the environment $q_{i}=\mathbf{x}_{i}\in\xi$, and the oracle’s answer $O(\mathbf{x}_{i})=z_{i}$ is typically given by collecting a new potentially noisy measurement $z_{i}\in\mathcal{F}$ at the location $\mathbf{x}_{i}$ based on the sensor readings. As summarized in Table 2, the performance metric $P$, i.e., evaluation metric, and thus acquisition function $A$, i.e., information criterion $I$, varies with the task. Reductions in the covariance matrix trace of a Gaussian process model or model uncertainty of an implicit neural representation are popular acquisition functions used in exploration and monitoring tasks. Other supervised approaches learn the acquisition function $A$ with deep neural networks (Hepp et al., 2018; Bai et al., 2017; Ly and Tsai, 2019) or directly predict the next best query $q_{i+1}^{*}$, e.g., a view pose to reconstruct a 3D object (Vasquez-Gomez et al., 2021; Mendoza et al., 2020).
Similarly, RL-based AIPP approaches presented in Sec. 4.2 rely on a reward function $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ based on the current and next state and an action $a_{i}\in\mathcal{A}$ to learn a policy that predicts the action $a_{i+1}^{*}\in\mathcal{A}$ maximizing the return, i.e., the sum of rewards. According to the definitions of the return in Eq. 3 and the AL problem in Eq. 5, the reward function in RL acts as the acquisition function in AL. In a POMDP describing the AIPP problem, the state $s$ can be seen as the current model $\theta$ trained on the history of collected queries $\{q_{1},\ldots,q_{i}\}$ and targets $\{O(q_{1}),\ldots,O(q_{i})\}$. The queries $q_{i}=a_{i}\in\mathcal{A}$ correspond to the previously chosen actions $a_{i}$, e.g., the next measurement location, and the oracle-generated targets $O(q_{i})=z_{i}\in\mathcal{O}$ correspond to the observations, e.g., sensor readings at measurement locations.
The AIPP approaches described so far assume static sensor models $p(z\,|\,F)$ to learn spatial environment models $\tilde{F}$ in Eq. 2. However, as environments $\xi$ are initially unknown, sensor models should be adapted to the robot environment to improve robot perception and, thus, the learned environment model $\tilde{F}$. This is crucial for continuous robot deployment in varying domains and for deep learning-based sensor models as they are known to transfer poorly to unseen scenarios.
To address this problem, a line of recent research considers using AIPP to actively gather data to improve a sensor model for robotic perception (Rückin et al., 2023, 2022b; Blum et al., 2019; Zurbrügg et al., 2022; Chaplot et al., 2021). In this problem setup, the information criterion $I$ in Eq. 1 is defined in terms of how potential future raw sensor readings, e.g., RGB images at certain environment locations, and their associated oracle-generated targets, e.g., semantic segmentation labels, impact the sensor performance. To solve this problem, the key idea is to link the AIPP objective to ideas from AL.
To date, this problem has only been studied in the context of pixel-wise semantic segmentation from RGB images (Rückin et al., 2023, 2022b; Blum et al., 2019; Zurbrügg et al., 2022; Chaplot et al., 2021). These approaches can be divided into two categories depending on how they implement the oracle for labeling new data: self-supervised methods exploit pseudo labels rendered from an online-built semantic map as an oracle, without considering a human annotator, while fully supervised methods only exploit dense pixel-wise human annotations as an oracle. In the first category, self-supervised methods can be used to fuse 2D semantic predictions from different viewpoints into a semantic 3D map (Zurbrügg et al., 2022; Chaplot et al., 2021). After a mapping mission is completed, the semantic 3D map is used to automatically render consistent 2D pseudo labels for re-training. In this line of work, Zurbrügg et al. (2022) propose an AIPP approach to plan viewpoints with high training data novelty for improved pseudo label quality. Similarly, Chaplot et al. (2021) train an RL agent to target low-confidence parts of a 3D map. Although these approaches do not require a human oracle Zurbrügg et al. (2022); Chaplot et al. (2021), they still rely on large labeled pre-training datasets to generate high-quality pseudo labels in new scenes. In contrast, fully supervised methods are more applicable to varying domains and environments. Blum et al. (2019) locally plan paths of high training data novelty. Rückin et al. (2022b) introduce a map-based planning framework supporting multiple acquisition functions (Rückin et al., 2023). However, these methods (Blum et al., 2019; Rückin et al., 2022b, 2023) require considerable human labeling effort to improve robotic vision.
These works highlight the versatility of applying AL methods to robotic learning through AIPP. The combination of AL and AIPP methods demonstrates a principled approach to integrating learning-based components and classically used uncertainty quantification techniques for AL. This integration represents a promising pathway towards robot deployment in more diverse environments with minimal human supervision. However, to harness the potential of AL and AIPP, future research must address multiple open problems. Existing uncertainty quantification methods for deep learning models are known to be overconfident (Abdar et al., 2021) and not well-calibrated in out-of-distribution scenarios. Further, fully and self-supervised approaches for sensor model improvement are limited in terms of human annotation effort and applicability to out-of-distribution environments, respectively. One possible direction is to take the strengths of both approaches and develop hybrid semi-supervised approaches (Rückin et al., 2024). We may be able to deploy more versatile robots by combining AIPP methods for environment monitoring and exploration (Sec. 4.1 to Sec. 4.3) with methods for targeted training data collection (Rückin et al., 2023, 2022b; Blum et al., 2019; Zurbrügg et al., 2022; Chaplot et al., 2021).
5 Applications
In this section, we categorize the AIPP works surveyed in our paper according to their target applications in robotics. Our goal is to assess the practical applicability of existing learning-based methods while pinpointing emerging trends and recognizing potential limitations.
Table 5 overviews AIPP applications in robotics and describes their relevant practical aspects. Our taxonomy considers four broad application areas: (i) environmental monitoring; (ii) exploration and search; (iii) semantic scene understanding; and (iv) active SLAM. For reference, in the last column, we also include the learning method: supervised learning (‘SL’); reinforcement learning (‘RL’); imitation learning (‘IL’); and/or active learning (‘AL’).
In general, our survey illustrates that the works are broadly applied in terrestrial, aquatic, and aerial domains and on diverse robot platforms. AIPP methods organically find usage in environmental monitoring scenarios to gather data about physical parameters, e.g., signal strength (Hollinger and Sukhatme, 2014; Denniston et al., 2023a; Bayerlein et al., 2021; Denniston et al., 2023b; Yang et al., 2023a; Song et al., 2015), temperature (Meliou et al., 2007; Rückin et al., 2022a; Westheider et al., 2023), or elevation (Viseras and Garcia, 2019; Zacchini et al., 2023). In particular, AIPP enables efficient data collection in large environments using sensors with limited coverage area, such as point-based sensors. For mapping, Gaussian processes are most commonly used to capture spatial correlations found in natural phenomena that are measured by a continuous variable. Several recent works use RL to learn the best data-gathering actions (Rückin et al., 2022a; Cao et al., 2023a; Denniston et al., 2023b), including achieving coordinated monitoring with multi-robot teams (Westheider et al., 2023; Zhu et al., 2023; Yanes Luis et al., 2023; Bayerlein et al., 2021).
AIPP strategies for exploration and search can be further broken down in terms of two distinct tasks: exploration of bounded unknown environments, i.e., outwards-facing scenarios, and active reconstruction of unknown objects or scenes, i.e., inwards-facing scenarios. The former necessarily accounts for possible obstacles in unknown regions and typically relies on depth-based sensing to construct volumetric maps as an environment is explored (Zwecher et al., 2022; Lodel et al., 2022; Choudhury et al., 2018; Tzes et al., 2023; Cao et al., 2023b). In contrast, most approaches for object reconstruction (Vasquez-Gomez et al., 2021; Lee et al., 2022; Jin et al., 2023; Sünderhauf et al., 2023; Mendoza et al., 2020) assume obstacle-free environments with a constrained hemispherical action space for a robot camera around the target object of interest. While NeRF-based methods are gaining popularity for this task (Jin et al., 2023; Sünderhauf et al., 2023; Lee et al., 2022; Yan et al., 2023; Zhang et al., 2023; Pan et al., 2022, 2023b), an open research challenge is applying them for exploration in more complex environments involving multiple objects and clutter.
Only a few works integrate semantic information or consider localization uncertainty in planning objectives for learning-based AIPP. Several approaches (Blum et al., 2019; Georgakis et al., 2022b; Zurbrügg et al., 2022; Rückin et al., 2023, 2022b, 2024) leverage semantics to improve deep learning-based sensor models via AL. Although semantic scene understanding provides valuable cues for learning complex AIPP behavior (Schmid et al., 2023), a key development barrier is the lack of high-quality semantic datasets relevant to the domains of robotic AIPP applications. The common assumption of perfect localization uncertainty represents a significant limitation to the applicability of existing methods, particularly in outdoor environments which may rely on imprecise Global Positioning System (GPS) signals (Caley et al., 2019; Hitz et al., 2017; Westheider et al., 2023; Choudhury et al., 2020; Hüttenrauch et al., 2019). Many approaches also neglect sensor uncertainty and noise, especially in environmental monitoring scenarios where models for specialized sensors are scarce. This is a major drawback since imperfect sensing has been shown to have significant effects on the quality of collected data and resulting planning strategy (Popović et al., 2020; Arora et al., 2019; Zacchini et al., 2023).
Our taxonomy reveals a gap in methods accounting for robot motion constraints while planning, with some methods assuming that the robot sensor can teleport between locations (Jin et al., 2023; Pan et al., 2022; Sünderhauf et al., 2023) or perform only incremental motions relative to its current position (Westheider et al., 2023; Viseras et al., 2021; Bayerlein et al., 2021). One open issue is the lack of guarantees on dynamic feasibility, which is essential for long-term path planning with agile, fast platforms such as small UAVs. Although there has been research on multi-robot missions (Westheider et al., 2023; Zhu et al., 2023; Yang et al., 2023a; Gao and Zhang, 2022; Hüttenrauch et al., 2019), there are several areas that have not been widely studied, such as tasks involving multi-sensor fusion (Ott et al., 2023), temporal changes in the environment (Caley et al., 2019; Marchant et al., 2014), and combining multiple problems simultaneously, e.g., semantic scene understanding and active SLAM.
Domain Robot platforms Sensors Obstacles? Planning space Sources of uncertainty Motion constraints Learning method
Environmental monitoring
(Velasco et al., 2020)
Air quality monitoring
UAV
Point-based gas sensor
2D
Mapping
(Choi and Cielniak, 2021)
Air quality monitoring
UGV
Point-based sensor
2D
Mapping
SL, RL
(Viseras and Garcia, 2019)
Elevation mapping
UAV team
Ultrasound height sensor
2D
Mapping
SL, RL
(Zeng et al., 2022)
Fruit monitoring
Robot arm
RGB-D camera
3D
Mapping
ROS $*{}^{*}$
RL
(Menon et al., 2023)
Fruit monitoring
Robot arm
RGB-D camera
3D
Mapping
ROS $*{}^{*}$
SL
(Denniston et al., 2023b)
Intensity monitoring
AUV/UGV/UAV
Generic camera, point-based sensor
2D/3D
Mapping
SL, RL
(Cao et al., 2023a)
Intensity monitoring
UGV
Point-based sensor
2D
Mapping
SL, RL
(Yang et al., 2023a)
Intensity monitoring
UGV team
Point-based sensor
2D
Mapping
SL, RL
(Song et al., 2015)
Intensity monitoring
UAV
Point-based sensor
2D
Mapping, sensing
Nonholonomic constraints
SL
(Hitz et al., 2017)
Lake bacteria monitoring
AUV
Point sensor for phycoerythrin fluorescence
3D
Mapping
Spline path
SL
(Duecker et al., 2021)
Ocean pollution monitoring
AUV
Pressure sensor, general water characterization
2D
Mapping
Spline path
RL
(Popović et al., 2020)
Precision agriculture
UAV
Generic camera
3D
Mapping, sensing, semantics
Spline path
SL
(Rückin et al., 2022a)
Precision agriculture
UAV
Generic camera
3D
Mapping, sensing
SL, RL
(Westheider et al., 2023)
Precision agriculture
UAV team
Generic camera
3D
Mapping, sensing
RL
(Arora et al., 2019)
Scientific data gathering
UGV
Generic camera, ultraviolet light source, laser scanner
2D
Mapping, sensing
Omnidirectional drive
RL
(Hollinger and Sukhatme, 2014)
Signal strength monitoring
AUV
Point-based sensor
2D
Mapping
SL
(Denniston et al., 2023a)
Signal strength monitoring
AUV/UGV
Generic camera, point-based sensor
2D
Mapping, path executability
RL
(Wei and Zheng, 2020)
Signal strength monitoring
UGV
Point-based sensor
2D
Mapping
SL, RL
(Caley et al., 2019)
Spatiotemporal field monitoring
AUV/UGV
Point-based sensor
2D
Mapping
SL
(Marchant et al., 2014)
Spatiotemporal field monitoring
UGV
Point-based sensor
2D
Mapping
Spline path
SL, RL
(Tzes et al., 2023)
Target tracking and localization
UGV team
RGB-D sensor
2D
Mapping, sensing
IL
(Meliou et al., 2007)
Temperature monitoring
Sensor network
Generic
2D
Mapping
SL
(Vivaldini et al., 2019)
Tree disease classification
UAV
RGB camera
2D
Mapping, semantics
SL
(Zacchini et al., 2023)
Underwater mapping
AUV
Sonar
3D
Mapping, sensing
Dubins vehicle model
SL
(Yanes Luis et al., 2023)
Water quality monitoring
AUV team
Water quality sensor
2D
Mapping
SL, RL
(Viseras et al., 2021)
Wildfire monitoring
UAV team
Thermal camera
2D
RL
(Bayerlein et al., 2021)
Wireless data harvesting
UAV team
Signal receptor
2D
Signal emissions
RL
Exploration and search
(Dai et al., 2022)
Camera view planning
UGV
RGB-D camera
3D
IL
(Yang et al., 2023b)
Exploration and landmark localization
UAV
Depth sensor
2D/3D
Mapping
RL
(Zhu et al., 2023)
Exploration and target search
UGV team
Depth sensor
2D
RL
(Choudhury et al., 2018)
Exploration, search-based planning
UGV/UAV
RGB-D camera
2D
Mapping
IL
(Lodel et al., 2022)
Floor plan exploration
UAV
Laser
2D
Mapping, sensing
Unicycle model
RL
(Liu et al., 2023)
Indoor exploration
UGV
Depth sensor
2D
RL, IL
(Zwecher et al., 2022)
Indoor exploration
UGV
Depth sensor
2D
Map prediction
SL, RL
(Bai et al., 2017)
Indoor exploration
UGV
Depth sensor
2D
Mapping, sensing
SL
(Tao et al., 2023)
Indoor exploration
UAV
Depth sensor
3D
Mapping, sensing
Spline path
SL
(Cao et al., 2023b)
Indoor exploration
UGV
Depth sensor
2D
Mapping, sensing
Local $*{}^{*}$ (Cao et al., 2022)
RL
  • 1.
    *** Motion constraints incorporated only during plan execution.
Table 5: Taxonomy of AIPP applications in robotics.
Domain Robot platforms Sensors Obstacles? Planning space Sources of uncertainty Motion constraints Learning method
Exploration and search
(Niroui et al., 2019)
Indoor exploration
UGV
Depth sensor
2D
Mapping, sensing
ROS $*{}^{*}$
RL
(Chen et al., 2019)
Indoor exploration
UGV
Depth sensor
2D/3D
Mapping, sensing
RL
(Shrestha et al., 2019)
Indoor exploration
UGV
Laser scanner
2D
Mapping
SL
(Li et al., 2023)
Indoor exploration
UGV
RGB-D camera
2D
Local planner using Fast Marching Method
SL
(Schmid et al., 2022)
Indoor exploration
UGV
RGB-D camera
2D
Mapping, sensing
SL
(Schmid et al., 2023)
Indoor exploration
UAV
RGB-D camera
3D
Mapping, sensing, scene completion
Velocity ramp model
SL
(Georgakis et al., 2022a)
Indoor exploration, point-goal navigation
UGV/UAV
Depth sensor
2D
Mapping, sensing
SL
(Ramakrishnan et al., 2020)
Indoor exploration
UGV/UAV
RGB-D camera
2D
Mapping, sensing
SL
(Gazani et al., 2023)
Infrastructure inspection
UAV
RGB-D camera
3D
RL
(Kumar et al., 2022)
Interactive exploration
Robot arm
RGB-D camera
3D
RL
(Chen et al., 2022)
Multi-object navigation
UGV
RGB-D camera
2D
Revolute joint camera
RL
(Novkovic et al., 2020)
Object finding
Robot arm
RGB-D camera
3D
Mapping, sensing
RL
(Best et al., 2019)
Object recognition, team orienteering
UGV team
2D Laser
2D
Mapping, sensing
RL
(Mendoza et al., 2020)
Object reconstruction
Any platform
Depth sensor
3D
SL
(Dhami et al., 2023)
Object reconstruction
Any platform
Depth sensor
3D
SL
(Vasquez-Gomez et al., 2021)
Object reconstruction
Any platform
Depth sensor
3D
Mapping
SL
(Pan et al., 2023b)
Object reconstruction
Any platform
RGB camera
3D
SL
(Pan et al., 2022)
Object reconstruction
Any platform
RGB camera
3D
Sensing
SL, AL
(Zhan et al., 2023)
Object reconstruction
Any platform
RGB camera
3D
Sensing
SL, AL
(Sünderhauf et al., 2023)
Object reconstruction
Any platform
RGB camera
3D
Sensing
SL, AL
(Jin et al., 2023)
Object reconstruction
Any platform
RGB camera
3D
Sensing
SL, AL
(Sukkar et al., 2019)
Object reconstruction
Robot arm
RGB-D camera
3D
Mapping, sensing
RL
(Lee et al., 2022)
Object reconstruction
Any platform
RGB-D camera
3D
Sensing
SL, AL
(Ran et al., 2022)
Object reconstruction
Any platform
RGB-D camera
3D
Sensing
SL, AL
(Yan et al., 2023)
Object reconstruction
Any platform
RGB-D camera
3D
Sensing
SL, AL
(Dhami et al., 2023)
Object reconstruction
UAV
Depth sensor
3D
ROS $*{}^{*}$
SL
(Pan et al., 2023a)
Object surface reconstruction
Any platform
RGB-D camera
3D
Sensing
ROS $*{}^{*}$
SL
(Cao et al., 2023c)
Persistent target monitoring
Robot team (UAV/UGV)
Generic camera
2D
Target position
SL, RL
(Ott et al., 2023)
Rover exploration
UGV
Drill, mass spectrometer
2D
Mapping, sensing
RL
(Choudhury et al., 2020)
Search and rescue
Any platform
Categorical survivor detection
2D
Mapping, sensing
RL
(Singh et al., 2009)
Search and rescue
Any platform
Optical sensor
2D
Sensing
RL
(Lim et al., 2016)
Search tasks, e.g., grasping pose or target search
Any platform
Binary measurements
3D
Sensing
(Reinhart et al., 2020)
Subterranean exploration
UAV
Depth sensor
3D
IL
(Qie et al., 2019)
Target assignment
UAV team
Not specified
2D
RL
(Gao and Zhang, 2022)
Target assignment
UAV team
Not specified
2D
Mapping, sensing
RL
(Hüttenrauch et al., 2019)
Target tracking
UAV team
Laser
2D
Mapping
Unicycle model
RL
(Ye et al., 2018)
Targeted object search
UGV
RGB camera
2D
Object recognition
RL
(Saroya et al., 2020)
Tunnel exploration
UGV
Depth sensor
2D
SL
(Ly and Tsai, 2019)
Urban exploration, surveillance
UGV/UAV
Depth sensor
2D
SL
(Hepp et al., 2018)
Urban exploration
UAV
Depth sensor
3D
Mapping, sensing
SL
(Zhang et al., 2023)
View planning for grasping
Any platform
RGB-D camera
3D
SL, AL
Domain Robot platforms Sensors Obstacles? Planning space Sources of uncertainty Motion constraints Learning method
Semantics
(Blum et al., 2019)
Land cover mapping
UAV
RGB camera
2D
Sensing
AL
(Georgakis et al., 2022b)
Indoor goal navigation
UGV
RGB-D camera
2D
Mapping, sensing
AL
(Zurbrügg et al., 2022)
Indoor semantic mapping
UGV
Laser, RGB-D camera
3D
Mapping, sensing, localization
AL
(Rückin et al., 2023)
Industrial and urban mapping, land cover analysis
UAV
RGB camera
2D
Mapping, sensing
(Rückin et al., 2022b)
Urban mapping
UAV
RGB camera
2D
Mapping, sensing
AL
(Rückin et al., 2024)
Urban mapping
UAV
RGB camera
2D
Mapping, sensing
AL
Active SLAM
(Hanlon et al., 2023)
Active visual localization
UGV
RGB-D camera
3D
Localization
SL
(Popović et al., 2020)
Indoor temperature monitoring
UGV/UAV
Laser, temperature sensor
2D/3D
Mapping, sensing, localization
SL
(Chen et al., 2020)
Exploration
UGV
Lidar
2D
Mapping, sensing, localization
RL
(Chen et al., 2021)
Exploration
UGV
Lidar
2D
Mapping, sensing, localization
RL
(Bartolomei et al., 2021)
Goal-based navigation
UAV
RGB-D camera
3D
Mapping, sensing, localization
Spline path
RL
6 Challenges and Future Directions
This section outlines the open challenges related to deploying learning-based AIPP in robotics settings and avenues for future work to overcome them.
6.1 Generalizability
Most of the AIPP methods surveyed were tested in the same or similar environments as those seen during training. The generalizability of deep learning-based systems is especially important because the neural networks that are often used in AIPP frameworks (Choudhury et al., 2018; Shrestha et al., 2019; Tao et al., 2023; Schmid et al., 2023; Li et al., 2020; Dai et al., 2022) are known to produce overconfident wrong predictions in out-of-distribution scenarios (Abdar et al., 2021). Future work should address these inherent generalization limitations to improve robustness. We identified one possible way to deal with deep learning-based vision uncertainties in Sec. 4.4 by combining uncertainty-aware deep learning and active learning approaches with AIPP (Chaplot et al., 2021; Zurbrügg et al., 2022; Rückin et al., 2023). Similarly, RL-based AIPP systems are increasingly popular (Cao et al., 2023a; Yang et al., 2023a; Cao et al., 2023b; Chen et al., 2021; Wei and Zheng, 2020; Choi and Cielniak, 2021). However, RL approaches degrade in performance when deployed in unseen simulated environments or real-world scenarios (Bousmalis et al., 2018; Zhao et al., 2020; Zhu et al., 2020). Future work on RL-based AIPP methods should focus on evaluating the learned policy in more diverse environments and in the real world. Further, new methods are needed to improve the generalization of RL for AIPP (Kirk et al., 2023). Other subfields of RL, such as meta-reinforcement learning (Gupta et al., 2018), could serve as inspiration for enabling efficient online policy adaption to unseen environments. Data augmentation and domain randomization during training could improve test-time policy robustness (Sadeghi and Levine, 2017).
6.2 Handling Localization Uncertainty
Effectively handling robot localization uncertainty is a key aspect in advancing learning-based AIPP in robotics. Localization uncertainties arise from various sources, including sensor noise, environmental changes, and the inherent limitations of localization technologies. However, most existing approaches assume perfect knowledge of the robot pose, both during planning and plan execution, which limits their practical applicability. Some works in active SLAM account for localization uncertainty via belief space planning (Chen et al., 2020, 2021) or modifying the planning objective (Popović et al., 2020) using graph-based representations of the robot state. Incorporating such probabilistic methods into learning-based AIPP methods and exploring transfer learning techniques (Chen et al., 2021; Genc et al., 2020) can enhance robustness.
6.3 Temporal Changes and Long-Term Planning
Addressing temporal changes and ensuring robust long-term planning in learning-based AIPP poses a critical challenge. Temporal changes, such as variations in environmental conditions, seasonal shifts, or the presence of dynamic obstacles, require AIPP methods to not only adapt to the immediate surroundings but also plan for sustained effectiveness over long durations. This involves developing strategies that consider the evolution of the environment and adjust the planned paths online. Whereas several works study spatiotemporal monitoring (Caley et al., 2019; Marchant et al., 2014; Cao et al., 2023c) and dynamic target tracking (Tzes et al., 2023; Hüttenrauch et al., 2019), there is a need for new approaches that can handle multiple moving objects in 3D environments. We also highlight the importance of developing such methods for agile platforms such as small UAV platforms while also accounting for their physical motion constraints, given that they often operate in dynamic and changing landscapes.
6.4 Heterogeneity
One challenge is the development of learning-based AIPP methods that exploit the benefits of heterogeneous robot teams or heterogeneous sensor modalities within a single robot. Leveraging heterogeneity in robotic systems has a high potential to improve sensing and monitoring performance. For instance, a tandem robotic system can combine the agility of a small UAV and the high-payload, detailed inspection capabilities of a UGV. A robot equipped with different sensors, e.g., a drill and a mass spectrometer (Ott et al., 2023), can measure different physical parameters during a single mission. To fully exploit these capabilities, we recommend investigating principled ways of incorporating heterogeneity, also considering collaboration aspects for effective decision-making.
6.5 Standardized Evaluation Methods
Another key challenge is the lack of standardized simulation environments, evaluation metrics, and benchmarks to assess new methods in AIPP. As highlighted in Sec. 3, most existing works rely on custom-made procedures and baselines for evaluation. While these reveal relative trends in performance, they do not ensure reproducible results across different approaches. Moreover, our survey catalogue [^1] https://dmar-bonn.github.io/aipp-survey/ indicates that very few works have publicly available datasets, documentation, and open-source code. In future work, it is essential to establish common evaluation pipelines to promote reproducibility and comparability. We recommend constructing such a framework in a modular fashion, allowing easy modification of various aspects, including the application scenario (robot platform, sensor setup, and environment) and algorithmic components (mapping method, planning objective, and planning algorithm). We believe such a development would also help guide future work towards more general approaches, rather than constrained application-specific methods, as well as foster stronger collaboration within the AIPP community.
7 Conclusion
In this survey paper, we analyzed the current research on AIPP in robotic applications, focusing especially on methods that exploit recent advances in robot learning. We introduced a unified mathematical framework for addressing problems in learning-based AIPP. Based on this formulation, we provided two complementary taxonomies from the perspectives of (i) learning algorithms and (ii) robotic applications. We provided a structured outline of the research landscape and highlighted synergies and trends. Finally, we discussed the remaining challenges for learning-based AIPP in robotics and outlined promising research directions.
8 Funding Sources
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2070 – 390732324.